# Easy-access online social media metrics can effectively identify misinformation sharing users

### v1

@@ Abstract
Misinformation poses a significant challenge studied extensively by researchers, yet acquiring data to identify primary sharers is costly and challenging. To address this, we propose a low-barrier approach to differentiate social media users who are more likely to share misinformation from those who are less likely. Leveraging insights from previous studies, we demonstrate that easy-access online social network metrics—average daily tweet count, and account age—can be leveraged to help identify potential low factuality content spreaders on X (previously known as Twitter). We find that higher tweet frequency is positively associated with low factuality in shared content, while account age is negatively associated with it. We also find that some of the effects, namely the effect of the number of accounts followed and the number of tweets produced, differ depending on the number of followers a user has. Our findings show that relying on these easy-access social network metrics could serve as a low-barrier approach for initial identification of users who are more likely to spread misinformation, and therefore contribute to combating misinformation effectively on social media platforms.
@@ Introduction
The problem of the global spread of misinformation has been recognised as one of the most pressing issues in today’s societies, prompting many researchers to dedicate efforts to understanding the issue as well as to developing potential solutions for tackling it [1]. Researchers have focused on a variety of areas to understand misinformation, examining the phenomenon from the perspectives of the misinformation producers, consumers, the messages themselves, and the contexts in which the messages are embedded [2]. While bots may play a significant role in spreading misinformation, the role of human users of sharing it remains a key element in the diffusion process [3]. Furthermore, as the spreading of misinformation carries the risk of subjecting people to manipulation and might influence behaviour in undesirable ways [4, 5], understanding the factors that affect individuals’ likelihood of sharing misinformation is essential.
@@
The reasons behind people’s tendency to share misinformation have been widely studied, and findings have shown that in many cases the decision of users to share misinformation does not stem from malicious intent, but it is explained by a failure to consider the accuracy of the piece of news shared [6]. Failure to consider accuracy when sharing content has been shown to have two main sources: lack of attention to accuracy when making the decision about sharing, or the inability to distinguish between true and false content [6]. Individuals who are – when asked – able to distinguish between true and false content, but consciously or unconsciously do not prioritise content accuracy when sharing, have been found to do so for the following reasons. First, when sharing content on social media, individuals are prone to prioritising social considerations, such as how sharing certain content affects their social status. Users might also use the content they share as a way to connect to others [7, 8, 9]. Second, it has been found that the high volume of information that individuals are required to process on social media results in less attention dedicated to each piece of content that the individual comes across [10]. Finally, the inability to distinguish between true and false content has been shown to be associated with characteristics such as limited cognitive skills [11], lack of analytical reasoning [12], and lower levels of digital literacy [13, 14].
@@
Interventions on social media have been shown to be successful across various contexts. Examples include prompting users to think about accuracy [11], and ‘inoculating’ users with small doses of misinformation and then refuting that piece of misinformation to increase awareness of how one can be manipulated into trusting content [15]. Given their success, it is compelling to explore ways in which social media users more likely to engage with misinformation sharing could be targeted with such interventions. While previous research presented above provides ample evidence on where to look for identifying individuals prone to misinformation sharing, obtaining data on these predictor characteristics can be complicated and/or expensive [16].
@@
In this paper, we set out to provide a low-barrier approach to identify users that are more likely to share misinformation, relying on insights from previous literature, and easy-access data from X (Twitter). Our aim in utilising X metrics as accessible proxies for the above outlined phenomena is to demonstrate their effectiveness as initial predictors of users’ susceptibility to misinformation sharing. We aim to show that with the use of basic information available on X accounts, it is possible to identify users who are more likely to share misinformation, providing a basis for low-cost interventions in the future.
For this purpose, we selected four metrics to test. Our criteria for selecting these metrics were their easy accessibility and their ability to serve as proxies for the significant determinants of misinformation sharing found in the literature. Our metric selection was informed by the literature to ensure efficiency, however we prioritised data accessibility instead of the selection of perfect proxies. Based on these criteria, we chose the following metrics to represent three lines of misinformation sharing determinants: follower count for social considerations, average daily tweet count and number of users followed for cognitive overload, and account age for digital literacy.
Our hypothesis for each metric is that it could serve as a useful indicator for distinguishing between low and high factuality users, with specific expectations for how each metric would affect user factuality. We expect that users with more followers and the ones who tweet frequently and follow more accounts are less factual, while users with a longer life on social media are more factual. Specifically, based on the literature on social considerations shifting attention away from accuracy we expect that users with a higher number of followers would have lower factuality scores than those with a lower number of followers [6]. Furthermore, we expect to find that the higher the number of tweets one shares and the higher the number of accounts one follows (i.e. the higher the information exposure), the less the attention each piece of content would receive, as suggested by [10]. Finally, we expect to find the set of high factuality users to be dominated by users with older accounts, and the set of low factuality users with newer accounts, based on work by [13].
@@ Results
Figure 1: Low factuality users tend to have higher follower count than high factuality users. Panel a: Comparing the distributions of follower count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followers compared to high factuality users (p-value $<$ 0.0001). The median value for low factuality users (orange dotted line) is 1635 followers, whereas for high factuality users (blue dotted line) it is 1155. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
@@
Figure 2: Low factuality users tend to have higher tweet count than high factuality users. Panel a: Comparing the distributions of tweet count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of tweets compared to high factuality users (p-value $<$ 0.0001). The median value for low factuality users (orange dotted line) is 2.99 tweets per day, whereas for high factuality users (blue dotted line) it is 1.78. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
Figure 3: Low factuality users tend to have higher followed account count than high factuality users. Panel a: Comparing the distributions of followed account count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followed accounts compared to high factuality users (p-value $<$ 0.0001). The median value for low factuality users (orange dotted line) is 2455 followed accounts, whereas for high factuality users (blue dotted line) it is 1914. The peak around 5000 is due to an X policy that limits the number of new followed accounts until the user obtains more followers. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
@@
Figure 4: Low factuality users tend to have lower number of days since registration than high factuality users. Panel a: Comparing the distributions of the number of days since registration between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly lower number of days compared to high factuality users (p-value $<$ 0.0001). The median value for low factuality users (orange dotted line) is 3134 days since registration, whereas for high factuality users (blue dotted line) it is 3493. Panel b: The MWU test score obtained from the empirical data (red dotted line) is lower than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
@@
Next, we employed multinomial regression to move from analysing the associations of factuality with individual social network features separately to considering the features all together. We do this in two stages. First, we perform multinomial regression with no interactions between the independent variables, second, we add interactions between explanatory variables. With conducting regression analysis, we are able to examine the effect of each key metric while simultaneously controlling for the effects of all other metrics. Thus, these models help us analyse the interplay of the metrics on factuality, and evaluate the extent to which these social network characteristics predict factuality scores.
@@
First, we consider and report the AMEs of the four social network metrics on factuality (Figure 5). Examining the AMEs of follower count on factuality group membership, we find that for both low and high factuality groups, holding other variables constant, the more followers the user has, the more likely they are to be in these groups, compared to the middle factuality group. This effect is slightly but significantly stronger for the low factuality group than for the high factuality group. However, as the effect of followers points towards the same direction for both low and high factuality groups compared to middle factuality, this characteristic appears to be a weak metric for distinguishing between low and high factuality users, and is rather fit to distinguish low and high factuality users from middle factuality ones.
@@
The number of users followed is negatively associated with both the high and low factuality groups, meaning that holding all other variables constant, following a high number of accounts makes a user more likely to be in the middle factuality category, while it makes them less likely to be in the low or high factuality categories. These findings are on the one hand consistent with our expectations, as we see a negative association between the number of users followed and high factuality, compared to middle factuality. Opposite to our expectation, we also see a negative association between the number of users followed and belonging to the low factuality category compared to belonging to the middle factuality category, however this effect is significantly weaker than the one belonging to high factuality. These effects show that there is a reverse U-shaped relationship between followed account count and the three factuality groups, both with low and high factuality users typically following fewer users than middle factuality users. These results suggest that the power of followed account count as a measure to classify low and high factuality users is therefore limited.
@@
The average number of tweets produced by the user per day is negatively associated with being in the high factuality group compared to the middle factuality group, i.e. holding all other variables constant, high factuality users tweet less on average. For the low factuality group, we find a positive association compared to the middle factuality group, i.e. holding all other variables constant, low factuality users tweet more on average. These results are in line with our hypothesis and previous results, confirming that average daily tweet count is a useful indicator in distinguishing between low and high factuality users.
@@
For days since registration, we see a positive effect for the high factuality, and the opposite for the low factuality group compared to the middle one. The results are therefore consistent with the outlined expectations and hypothesis, indicating that, holding all other variables constant, longer presence on X is associated with higher factuality, and that the metric is useful in distinguishing between low and high factuality users.
@@
Figure 5: Average marginal effects of social network characteristics on factuality. All examined social network metrics are significantly associated with factuality, and the differences in the effects between low factuality (marked with orange) and high factuality (marked with blue) users is significant in all four cases. Zero effect would mean users are equally likely to be low or high factuality as middle factuality. Tweet count: higher tweet count means the user is more likely to be low factuality, and less likely to be high factuality than middle factual. Followed account count: higher followed account count means the user is less likely to be low factuality, as well as to be high factuality, but the latter effect is stronger. Follower count: higher follower count means the user is more likely to be low factuality, as well as to be high factuality, with the former effect being somewhat stronger. Days Since Registration: higher number of days means the user is more likely to be high factuality, and less likely to be low factuality.
@@
Finally, we examined the interactions between the effects of the four social network metrics on factuality to unfold potential interplay of the independent variables. We have found that some of the social network metrics interact significantly with one another on factuality. To illustrate the interactions we divided the users into two equally sized groups based on the given social network metrics, e.g. follower count, and plotted the effects at the median of each group.
Our analysis reveals differences regarding the effect of the number of users followed on factuality (Figure 6, panel a). The effect of followed account count on high factuality is strongly negative among those with smaller follower base (less popular users), while slightly positive among those with larger follower base (popular users). In other words, less popular users are less likely to exhibit high factuality as they follow more users; whereas popular users show a slightly increased likelihood of high factuality with more followed accounts. This shows that the effect depends heavily on the number of followers the user has. Furthermore, this result shows that this effect is different from the main effect of followed account count, as the main effect showed a negative relationship, but in the case of popular users we found a slightly positive one. We can observe that among those who follow a low number of accounts, less popular users are more likely to be highly factual; while among those who follow many accounts, popular users are more likely to be highly factual. For the low factuality group similar trends are observed for both categories of follower count, where both less popular and popular users are less likely to exhibit low factuality as they increase the number of users they follow, with somewhat stronger effect for popular users.
In relation to average daily tweet count, our analysis (Figure 6, panel b) shows that as users increase their tweet activity, they are less likely to belong to the high factuality category and more likely to belong to the low factuality category, regardless of their number of followers. However, we do find differences in the strength of the effect. In the high factuality group the effect is more pronounced among less popular users, compared to the effect observed among popular users. Among those with a low number of average daily tweets, less popular users are more likely to be highly factual; while among those who tweet a lot, more popular users are more likely to be highly factual. Conversely, in the low factuality group, the relationship is less pronounced among less popular users. There we can observe that among those who only post a few tweets on average there is no difference between less popular and popular users; however, among those who tweet a high amount, popular users are more likely to belong to the low factuality group than less popular users.
@@
In examining the interactions of the effects among the rest of the explanatory variables we found only slight differences. In none of the other cases was there a difference in the direction of the effect based on another variable. The observed differences are small, regarding either the magnitude or the strength of the effects (see the full set of explanatory variable interactions in Supplementary Figure 10).
@@
Figure 6: Average marginal effect interactions between follower count and followed account count and average daily tweet count. Panel a: The effect of the number of accounts followed on factuality differs between users with few and many followers. The effect is opposite for the high factuality group, and different in strength for the low factuality group. Panel b: The effect of the average daily tweet count on factuality differs between users with few and many followers. The effect is opposite for the high factuality group, and different in strength for the low factuality group.

### v4

# Easy-access online social media metrics are associated with misinformation sharing activity

@@ Abstract
Misinformation poses a significant challenge studied extensively by researchers, yet acquiring data to identify primary sharers is time-consuming and challenging. To address this, we propose a low-barrier approach to differentiate social media users who are more likely to share misinformation from those who are less likely. Leveraging insights from previous studies, we demonstrate that easy-access online social network metrics—average daily tweet count, and account age—can be leveraged to help identify potential low factuality content spreaders on X (previously known as Twitter). We find that higher tweet frequency is positively associated with low factuality in shared content, while account age is negatively associated with it. We also find that some of the effects differ depending on the number of accounts a user follows. Our findings show that relying on these easy-access social network metrics could serve as a low-barrier approach for initial identification of users who are more likely to spread misinformation, and therefore contribute to combating misinformation effectively on social media platforms.
@@ Introduction
The problem of the global spread of misinformation has been recognised as one of the most pressing issues in today’s societies, prompting many researchers to dedicate efforts to understanding the issue as well as to developing potential solutions for tackling it [1]. Researchers have focused on a variety of areas to understand misinformation, examining the phenomenon from the perspectives of the misinformation producers, consumers, the messages themselves, and the contexts in which the messages are embedded [2]. While bots may play a significant role in spreading misinformation, the role of human users in sharing it remains a key element in the diffusion process [3]. Furthermore, as the spreading of misinformation carries the risk of subjecting people to manipulation and might influence behaviour in undesirable ways [4, 5], understanding the factors that affect individuals’ likelihood of sharing misinformation is essential.
@@
The reasons behind people’s tendency to share misinformation have been widely studied, and findings have shown that in many cases the decision of users to share misinformation does not stem from malicious intent, but rather reflects a failure to consider the accuracy of the piece of news shared [6]. Two pathways have been proposed for this failure: limited attention to accuracy in the moment of sharing, or difficulty in distinguishing between true and false content [6]. In particular, even individuals who can distinguish between true and false information more easily often fail to apply this discernment at the point of sharing, due to contextual or cognitive factors. First, when sharing content on social media, individuals are prone to prioritising social considerations, such as how sharing certain content affects their social status. Users might also use the content they share as a way to connect to others [7, 8, 9]. Second, limited attention caused by high information volume on social media platforms reduces users’ ability to reflect on the accuracy of each piece of content they encounter [10]. These processes have been linked to various individual-level traits, including limited cognitive skills [11], lack of analytical reasoning [12], and lower levels of digital literacy [13, 14].
@@
Interventions on social media have been shown to be successful across various contexts. Examples include prompting users to think about accuracy [11], and ‘inoculating’ users with small doses of misinformation and then refuting that piece of misinformation to increase awareness of how one can be manipulated into trusting content [15]. Given their success, it is compelling to explore ways in which social media users who are more likely to engage with misinformation sharing could be targeted with such interventions. While previous research presented above provides ample evidence on where to look for identifying individuals prone to misinformation sharing, obtaining data on these predictor characteristics can be complicated, time-consuming, and/or expensive [16].
@@
In this paper, we set out to provide a low-barrier, explanatory approach to assess whether simple, easily accessible account-level characteristics are systematically associated with users’ likelihood of sharing misinformation, drawing on insights from previous literature and using easy-access data from X (Twitter). Our aim in utilising X metrics as accessible proxies for the above outlined phenomena is to explore whether they can serve as early-warning signals of lower factuality tendencies among users. We aim to show that basic information available on X accounts can help surface broad patterns associated with lower factuality, offering a potential foundation for scalable and low-cost interventions.
For this purpose, we selected four metrics to test: (1) follower count, (2) average daily tweet count, (3) followed account count, and (4) account age. In our criteria for selecting these metrics we prioritised metrics that are widely available across platforms; not reliant on user self-report, thereby reducing biases associated with self-disclosure; and retrievable in a structured and standardized way, facilitating potential cross-platform generalizability; as well as the metrics’ potential ability to serve as proxies for the significant determinants of misinformation sharing found in the literature.
While we focus on these metrics due to their accessibility, we recognize that other account-level characteristics may also be useful for understanding users’ likelihood of sharing low-factuality content. For example, features such as whether a user provides a location or the complexity of their bio text could add additional explanatory power. Additionally, we acknowledge that alternative methods—such as timeline content or network-based analysis—may yield greater accuracy in detecting misinformation [17]. However, these approaches typically require substantially more data collection and processing effort, or more advanced computation. In contrast, the metrics we rely on—follower and following counts, tweet frequency, and account age—are embedded in the basic profile metadata and can be obtained and analyzed with minimal technical and infrastructural requirements. This makes them especially valuable in contexts where data access is restricted, resources are limited, or scalability is a priority. In other words, our metric selection was informed by the literature to ensure efficiency, however we prioritised data accessibility instead of the selection of perfect proxies.
Based on these criteria, we chose the following metrics to represent three lines of misinformation sharing determinants: follower count for social considerations, average daily tweet count and number of users followed for cognitive overload, and account age for digital literacy. Our hypothesis for each metric is that it could serve as a useful indicator for distinguishing between low and high factuality users, with specific expectations for how each metric would affect user factuality. We expect to find users with more followers and the ones who tweet frequently and follow more accounts to be less factual, while users with older accounts are expected to be more factual. The expectation regarding follower count is based on research suggesting that social motivations—such as maintaining reputation or appealing to an audience—can divert attention from content accuracy [6]. A higher number of followers may increase perceived audience pressure or reputational concerns, which have been shown to influence sharing decisions. Similarly, we hypothesize that users who produce large volumes of tweets or follow many accounts experience a form of cognitive strain or information overload, reducing the attention they can dedicate to assessing accuracy [10, 18]. We acknowledge that platform algorithms shape users’ content exposure, and that a user’s feed may include messages from both followed and non-followed accounts. Nonetheless, following more accounts likely increases the potential content diversity and volume a user may encounter, even within algorithmically filtered feeds. Finally, we interpret account longevity as a rough proxy for digital experience or literacy. More precisely, we treat account age as a behavioral approximation of platform familiarity and accumulated exposure, rather than digital literacy in a strict sense. We also note that account age may correlate with user age or other demographic traits, which complicates its interpretation. Account longevity could reflect a mix of factors — including but not limited to digital literacy — such as habitual platform use, early adopter behavior (which has been linked to higher media savvy [19]), and greater exposure to platform norms and misinformation corrections. This interpretation aligns with prior work on digital behavior and literacy [20], while also acknowledging the conceptual limitations of using account age as a stand-in for digital skills. Hence, although several techniques exist to infer users’ propensity to consume unreliable sources, the development of a method based on easily accessible metrics—grounded in the literature on misinformation susceptibility—may offer a useful framework for understanding broad patterns of vulnerability and the design of targeted countermeasures to curb the spread of false or inaccurate information.
@@ Results
To ensure that our analysis focuses on regular users rather than automated or public accounts (e.g., news sites, celebrities), we applied specific filtering thresholds. In our main analysis, we used the following middle-level filtering criteria: only non-verified accounts were included; accounts were limited to a maximum of 32 tweets per day on average; follower count was capped at 10,000; follower-to-following ratio had to be 5 or lower; followed account count was capped at 10,000; following-to-follower ratio had to be 10 or lower. We report these filtering criteria and their justification in Filtering the dataset for regular users and Table 1.
To construct low and high factuality groups, we divided users into factuality tertiles based on their domain-sharing profiles—focusing on the bottom and top 30% in the main analysis, and testing 25% and 35% thresholds for robustness checks (see Supplementary Information).
@@
Figure 1: Low factuality users tend to have higher follower count than high factuality users. Panel a: Comparing the distributions of follower count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followers compared to high factuality users (p-value << 0.0001). The median value for low factuality users (orange dotted line) is 1642 followers, whereas for high factuality users (blue dotted line) it is 1125. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
@@
Figure 2: Low factuality users tend to have higher tweet count than high factuality users. Panel a: Comparing the distributions of tweet count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of tweets compared to high factuality users (p-value << 0.0001). The median value for low factuality users (orange dotted line) is 3.26 tweets per day, whereas for high factuality users (blue dotted line) it is 1.94. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
Figure 3: Low factuality users tend to have higher followed account count than high factuality users. Panel a: Comparing the distributions of followed account count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followed accounts compared to high factuality users (p-value << 0.0001). The median value for low factuality users (orange dotted line) is 2415 followed accounts, whereas for high factuality users (blue dotted line) it is 1845. The peak around 5000 is due to an X policy that limits the number of new followed accounts until the user obtains more followers. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
@@
Figure 4: Low factuality users tend to have lower number of days since registration than high factuality users. Panel a: Comparing the distributions of the number of days since registration between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly lower number of days compared to high factuality users (p-value << 0.0001). The median value for low factuality users (orange dotted line) is 3242 days since registration, whereas for high factuality users (blue dotted line) it is 3578. Panel b: The MWU test score obtained from the empirical data (red dotted line) is lower than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
@@
Next, we employed multinomial regression to move from analysing the associations of factuality with individual social network features separately to considering the features all together. We do this in two stages. First, we perform multinomial regression with no interactions between the independent variables, second, we add interactions between explanatory variables. With conducting regression analysis, we are able to examine the effect of each key metric while simultaneously controlling for the effects of all other metrics. Thus, these models help us analyse the interplay of the metrics on factuality, and evaluate the extent to which these social network characteristics are associated with factuality scores.
@@
First, we consider and report the AMEs of the four social network metrics on factuality (Figure 5). Examining the AMEs of follower count on factuality group membership, we find that the more followers users have, the more likely they are to be in the low factuality group and less likely to be in the high factuality group, compared to the middle factuality group. Specifically, a one standard deviation increase in follower count is associated with a 1.3 percentage point increase on average in the probability of belonging to the low factuality group, and on average a 2.3 percentage point decrease in the probability of belonging to the high factuality group, compared to the middle factuality group. This suggests that, even when accounting for other variables, follower count remains a useful indicator for distinguishing between high and low factuality users.
@@
The number of users followed is negatively associated with both the high and low factuality groups, meaning that holding all other variables constant, following a high number of accounts makes a user more likely to be in the middle factuality category, while it makes them less likely to be in the low or high factuality categories. Quantitatively, a one standard deviation increase in the number of followed accounts is associated with a 1.0 percentage point decrease on average in the probability of being in the low factuality group and on average a 1.9 percentage point decrease in the probability of being in the high factuality group, compared to the middle factuality group. These findings are on the one hand consistent with our expectations, as we see a negative association between the number of users followed and high factuality, compared to middle factuality. Opposite to our expectation, we also see a negative association between the number of users followed and belonging to the low factuality category compared to belonging to the middle factuality category, however this effect is significantly weaker than the one belonging to high factuality. These effects show that there is a reverse U-shaped relationship between followed account count and the three factuality groups, both with low and high factuality users typically following fewer users than middle factuality users. These results suggest that the power of followed account count as a measure to classify low and high factuality users is therefore limited.
@@
The average number of tweets produced by the user per day is negatively associated with being in the high factuality group compared to the middle factuality group, i.e. holding all other variables constant, high factuality users tweet less on average. For the low factuality group, we find a positive association compared to the middle factuality group, i.e. holding all other variables constant, low factuality users tweet more on average. More specifically, a one standard deviation increase in average daily tweet count is associated with a 0.23 percentage point increase on average in the probability of being in the low factuality group, and on average a 1.2 percentage point decrease in the probability of being in the high factuality group, compared to the middle factuality group. These results are in line with our hypothesis and previous results, confirming that average daily tweet count is a useful indicator in distinguishing between low and high factuality users.
@@
For days since registration, we see a positive effect for the high factuality, and the opposite for the low factuality group compared to the middle one. Numerically, a one standard deviation increase in days since registration is associated with a 2.2 percentage point decrease on average in the probability of being in the low factuality group, and on average a 1.5 percentage point increase in the probability of being in the high factuality group, compared to the middle factuality group. The results are therefore consistent with the outlined expectations and hypothesis, indicating that, holding all other variables constant, longer presence on X is associated with higher factuality, and that the metric is useful in distinguishing between low and high factuality users.
To assess the robustness of our results, we tested the consistency of the observed effects across eight alternative dataset constructions, varying both the user filtering criteria and the factuality thresholding. We found that tweet frequency and days since registration consistently distinguished between high and low factuality users across all variations, with stable effect directions and magnitudes. Follower count also showed a qualitatively consistent pattern, with a stable difference between high and low factuality groups, although the effect size varied somewhat across conditions. The number of followed accounts showed more variation, but the distinction between high and low factuality users—particularly the stronger negative association for high factuality users—was preserved across nearly all dataset versions. These results reinforce the general robustness of our findings while also highlighting that some metrics (e.g., follower count and followed account count) are more sensitive to dataset construction choices than others (see Figure 10 in the Supplementary Information).
@@
Figure 5: Average marginal effects of social network characteristics on factuality. All examined social network metrics are significantly associated with factuality, and the differences in the effects between low factuality (marked with orange) and high factuality (marked with blue) users is significant in all four cases. Zero effect would mean users are equally likely to be low or high factuality as middle factuality. Tweet count: higher tweet count means the user is more likely to be low factuality, and less likely to be high factuality than middle factual. Followed account count: higher followed account count means the user is less likely to be low factuality, as well as to be high factuality, but the latter effect is stronger. Follower count: higher follower count means the user is more likely to be low factuality, and less likely to be high factuality. Days Since Registration: higher number of days means the user is more likely to be high factuality, and less likely to be low factuality.
@@
Finally, we examined the interactions between the effects of the four social network metrics on factuality to unfold potential interplay of the independent variables. We have found that some of the social network metrics interact significantly with one another on factuality. To illustrate the interactions we divided the users into two equally sized groups based on the given social network metrics, e.g. followed account count, and plotted the effects at the median of each group.
Our analysis reveals differences regarding the effect of the average daily tweet count on factuality (Figure 6, panel a). In examining average daily tweet count across different levels of followed account count and its association with factuality levels we find that among those with few accounts followed the average tweet count does not affect the likelihood of belonging to the high factuality group. However, an increase in tweet activity is associated with an increased likelihood of belonging to the high factuality group among those who follow more accounts. For the low factuality group, tweet count is negatively associated with low factuality and this effect is stronger among those who follow more accounts. There is no significant difference in the likelihood of belonging to the low factuality group among those who tweet infrequently, however, among those who tweet frequently, users following few accounts are way more likely to have low factuality than users who follow more accounts.
In relation to days since registration, our analysis (Figure 6, panel b) shows that among users who follow fewer accounts, those who have been registered on the platform for a longer duration are more likely to be in the high factuality category, while among those who follow more accounts, the trend is the opposite: the longer they are on the platform, the less likely that they belong to the high factuality group. The relationship between days since registration and low factuality does not vary based on followed account count. The relationship is consistently negative between days since registration and low factuality, indicating that longer registration periods are associated with a decreased likelihood of low factuality, however, this effect is somewhat more pronounced among users who follow fewer accounts.
@@
In examining the interactions of the effects among the rest of the explanatory variables we found only slight differences. In only one of the other cases was there a difference in the direction of the effect based on another variable (effects of followed account on high factuality by follower count). The rest of the observed differences are small, regarding either the magnitude or the strength of the effects (see the full set of explanatory variable interactions in Supplementary Figure 11).
@@
Figure 6: Average marginal effect interactions between followed account count, average daily tweet count, and days since registration. Panel a: The effect of tweet activity on factuality differs by followed account count, with positive effects on high factuality and negative effects on low factuality primarily among those who follow many accounts. Panel b: The effect of days since registration on factuality also varies by followed account count, with opposite trends for high factuality and consistent negative association with low factuality.